OpenStudy (anonymous):

-x squared-3x+4=0 solve by completing the square.

OpenStudy (anonymous):

-x^2 - 3x + 4 = 0 x^2 + 3x - 4 = 0 x^2 + 3x + 9/4 - 4 - 9/4 = 0 (x + 3/2)^2 - 25/4 = 0 (x + 3/2)^2 = 25/4 x + 3/2 = +/- 5/2 x = -4 or +1